# GD doesn’t make the cut: Three ways that non-differentiability affects neural network training

### v2

@@ 3 NDGM and the LASSO penalty
Zero weights play a critical role in model compression (Blalock et al., 2020; Han et al., 2015a, b; Li et al., 2016), and it is desirable to force weights down to zero whenever we can do so without sacrificing too much accuracy. A common approach to induce sparsity is to penalize the $L_{1}$ norm of the model weights, with the penalizing factor commonly referred to as the LASSO penalty. In the statistics literature, loss functions involving penalized $L_{1}$ norms are typically minimized using specialized algorithms designed for different use cases (Efron et al., 2004; Friedman et al., 2008; Mazumder & Hastie, 2012; Tibshirani, 1996; Tibshirani et al., 2005; Zou & Hastie, 2005). However when training neural networks, the penalized $L_{1}$ norm is used with arbitrary non-convex functions, with the understanding that running NDGM on the problem results in a near sparse optimal solution (Bengio, 2012; Goodfellow et al., 2016). As stated in Scardapane et al. (2017), ”Even if (the LASSO) has a non differentiable point in 0, in practice this rarely causes problems to standard first- order optimizers. In fact, it is common to simultaneously impose both weight-level sparsity with the $L_{1}$ norm, and weight minimization using the $L_{2}$ norm, resulting in the so-called elastic net penalization”. This expectation is not being met in reality, and there is strong evidence of users struggling with the interpretation of their LASSO results in the PyTorch forums  [^3] https://rb.gy/kh0qr  [^4] https://rb.gy/2vsry  [^5] https://rb.gy/6dfli. Given the wide variety of critical applications in which the LASSO finds use (Ghosh et al. (2021); Li et al. (2011); Li & Sillanpää (2012); Ogutu et al. (2012) to state a few), it is critical to understand the reliability of the NDGM solution for the LASSO problem.
@@
The NDGM iteration for (7) is described by the recursion
$\beta_{N}[k]=\beta_{0}[k]+(N-2m)\alpha\lambda_{1}=0,$ (9)
Proposition 3.2.
The sequence $\{\beta_{t}\}_{t=0}^{\infty}$ described in the previous proposition does not converge. Furthermore, there is an integer $N_{0}$ such that for every $N\geq N_{0}$, the sequence oscillates between two vectors : $\gamma_{1}$ and $\gamma_{2}$, where $||\gamma_{1}||_{\infty}<\alpha\lambda_{1}$, and $||\gamma_{2}||_{\infty}<\alpha\lambda_{1}$.
Proof.
We prove the results assuming $\beta_{0}[k]>0$; the proof is similar when $\beta_{0}[k]<0$. From Proposition 3.1, we know that for any $N>0$, $\beta_{N}[k]\neq 0$ with probability 1. Therefore, (8) implies that starting from any $\beta_{0}[k]>0$, the sequence will decrease monotonically till it reaches a value between 00 and $\alpha\lambda_{1}$. Accordingly, let $\beta_{n_{k}}[k]=\gamma_{k}$ for some $n_{k}\geq 0$, with $\gamma_{k}\in(0,\alpha\lambda_{1})$. Since $\gamma_{k}-\alpha\lambda_{1}<0$, completing the recursion in (8) gives
Equation (10) suggests the surprising result that with the same learning rate and initialization, a larger value of $\lambda_{1}$ can result in a NDGM solution with a larger $L_{1}$ norm, which completely defeats the purpose of the LASSO penalty. To better understand the issue, consider the 2D case, with the weights initialized as $\beta_{0}=[0.5053,0.5053]$. With $\lambda_{1}=1$ and $\alpha=0.01$, Proposition 3.2 states that the NDGM sequence will oscillate between $[0.0053,0.0053]$ and $[-0.0047,-0.0047]$, but with $\lambda_{1}=100$, and the same learning rate and starting point, the solution will oscillate between $[0.5053,0.5053]$ and $[-0.4947,-0.4947]$. Thus, at the point when NDGM appears to converge, the solution with $\lambda_{1}=1$ will have an $L_{1}$ norm of at most 0.0106, but solution with $\lambda_{1}=100$ case will have an $L_{1}$ norm of atleast 0.9894. Furthermore, note that reducing the step-size does not resolve this paradox. If, in both cases, we decrease the learning rate to $\alpha=0.001$ once the algorithm seems to have converged, then with $\lambda_{1}=1$, the NDGD sequence will oscillate between $[0.0003,0.0003]$ and $[-0.0007,-0.0007]$. However, with $\lambda_{1}=100$, the NDGM sequence will oscillate between $[0.0053,0.0053]$ and $[-0.0947,-0.0947]$, and the paradox persists.
To demonstrate the same issue in the broader setting, consider the loss function for the general LASSO problem given by
Connections to the subgradient method:
The counter-intuitive behavior seen in the example above can be explained using the subgradient method. Since the loss function described in equation (7) satisfies the Lipschitz condition (see equation (6)) with $L=\lambda_{1}$, we have from Boyd et al. (2003) (Section 2 on the Constant Step Size) that
@@
Therefore, for the same value of $\alpha$, the larger value of $\lambda_{1}$ will result in a larger error in estimation, which is precisely what we are seeing above.
Diminishing v/s reducing step sizes:
This illustration also highlights the contrast between the diminishing and reducing step size regimes outlined in the preliminaries. Under the diminishing step-size regime, $\alpha\rightarrow 0$ in the theoretical limit, and therefore, Proposition 3.1 guarantees that the optimal solution for (7) will be 0 for all $\lambda_{1}>0$. On the other hand, in the reducing step-size regime, $\alpha>0$ at the end of training, and therefore, the error in estimation becomes a critical determinant of the optimal value.
@@
In their seminal work on network pruning, Han et al. (2015b) propose the utilization of $L_{1}$ regularization to encourage weights to approach 0, facilitating the pruning process. This approach is widely adopted in the literature (Wen et al., 2016; Xiao et al., 2019; Yang et al., 2019), and is explicitly endorsed in NVIDIA’s guides for training standard object detection models such as RetinaNet and YOLOv4666See the recommendation of the regularizer in the Training config section of https://rb.gy/20y13m and https://rb.gy/i1sb6s. To challenge the general applicability of this suggestion, we fine-tune two pretrained VGG16 models using the $L_{1}$-penalized cross entropy loss with a learning rate of $\alpha=0.01$, with one model using a regularization parameter of $\lambda_{1}=0.001$ and the other with $\lambda_{1}=0.1$. At the end of fine-tuning, the $L_{1}$ norm of the weights in the model fine-tuned with $\lambda_{1}=0.1$ is more than 30 times larger (69178.44) than that of the model fine-tuned with $\lambda_{1}=0.001$ (2113.83). Furthermore, the latter model has 99.7% of its weights less than $10^{-5}$, whereas the former model has only 0.01% of its weights less than $10^{-5}$, further underscoring the discrepency. We conducted experiments with different values of $\alpha$, and the widely used Adam optimizer (Kingma & Ba, 2014), and in each case, we observed the same paradox. Our code is provided in the appendices.
@@
While NDGMs do not inherently generate sparse solutions, iterative methods have been developed to produce sparse solutions for the $L_{1}$ regularized linear problem (see Table 1 in Bottou (2010)), and $L_{1}$ regularized neural networks in general (Ma et al., 2019; Siegel et al., 2020). The realization that NDGMs lack the capability to produce sparse solutions predates the widespread adoption of deep learning. As mentioned in (Xiao, 2009), “… stochastic gradient descent (SGD), has limited capability of exploiting problem structure in solving regularized learning problems”. Despite this, a prevalent misconception persists in the deep learning literature, erroneously suggesting that such algorithms yield sparse solutions, and that the solutions generated by deep learning frameworks like PyTorch (Paszke et al., 2017) ought to be comparable to those generated using traditional machines learning frameworks like glmnet (Friedman et al., 2021) or scikit-learn (Kramer & Kramer, 2016). We point out that there exits a fundamental disparity between traditional machine learning frameworks like glmnet and deep learning frameworks like PyTorch. glmnet focuses on a narrow set of optimization problems, concealing implementation intricacies while guaranteeing optimality. Conversely, PyTorch provides greater flexibility in constructing models but lacks similar guarantees, placing the responsibility for solution quality on users. Attaining a solution at a non-differentiable point poses a challenging problem, underscoring the significance of exploring related studies and tools rather than opting for a simplistic implementation, particularly when precision is paramount.
Connections to the Edge of Stability:
@@ 4 NDGM and the Edge of Stability
In recent years, there has been a growing interest on the topic of “unstable convergence”. The literature on the subject (Ahn et al., 2022a, b; Arora et al., 2022; Chen & Bruna, 2022; Cohen et al., 2021; Li et al., 2022) is motivated by the finding that unlike convex quadratic forms, gradient descent on neural networks do not diverge to $\infty$ even when the learning rate is greater than $\alpha^{*}=2/\eta$, where $\eta$ is the dominant eigenvalue of the loss function’s Hessian. Instead, the loss function has been empirically demonstrated to converge “unstably”, reducing non-montonically on the long run. $\alpha=\alpha^{*}$ is referred to in the literature as the ”Edge of Stability”, since it separates regions of ”stable convergence” ($\alpha<\alpha^{*}$) from regions of “unstable convergence” [^7] We define unstable convergence to mean “non-divergence to $\infty$”. Our definition is adapted from Definition 1.1 of Arora et al. (2022). However, we deviate from their definition due to its dependence on the existence of the Hessian at all points between the current and next iterate, a condition that need not be satisfied in non-smooth neural networks. ($\alpha>\alpha^{*}$).
The derivation of the Edge of Stability condition assumes that the loss function, $f$, is $L-smooth$ (see equation (5)). Since the result has been demonstrated in most convex optimization settings (Jacot et al., 2018; Li & Liang, 2018), it is conjectured that the result ought to hold even in the case of non-convex, non-differentiable neural networks. This conjecture is incorrect even for Lipschitz continuous, differentiable, convex functions satisfying (6) as shown in the proposition below.
@@
In the previous section, we analytically derive the unstable convergence for the LASSO problem described in (7). An example of a differentiable non-smooth convex function satisfying Proposition 4.1 is the Huber loss function, which finds use in many state of the art object detection models (Girshick, 2015; Liu et al., 2016; Ren et al., 2015). The Huber loss function is non-smooth because it is once but not twice differentiable. For a regression problem involving an arbitrary $50\times 1$ response vector, $\mathbf{y}=[y_{1},y_{2}\ldots y_{N}]^{T}$, and an arbitrary $50\times 200$ data matrix, $\mathbf{Z}$, the Huber loss function is defined as
@@
where $z_{i}$ is the $i^{th}$ column of $\mathbf{Z}$. Figure 3 shows that the NDGM sequence for (13) does not diverge toward $\infty$, even with a high learning rate, such as $\alpha=10$, which aligns with our expectations. Furthermore, the figure reveals an intriguing loss pattern when $\alpha=1$: the loss initially increases before the oscillations become noticeable. This peculiar behavior has been documented in the literature (Ahn et al., 2022b; Arora et al., 2022), although a conclusive explanation for this phenomenon remains elusive at this time.
Proposition 4.1 suggest that it is unclear and the examples that follow show that the common approach of extrapolating results developed for $L-smooth$ loss functions to the non-smooth or non-differentiable case is problematic. In the next subsection, we demonstrate the problems with this approach by providing an example of an extrapolation which breaks down in the non-smooth case.
Subquadratic losses and unstable convergence
In section 3.1 of Ma et al. (2022), the authors employ second-order finite differences to investigate the curvature of VGG16 and ResNet in the vicinity of the local minimum. They conclude that subquadratic growth of the loss function around this minimum point explains the edge of stability phenomenon. However, their analysis is flawed because VGG16 and ResNet, as ReLU networks, have local minima that are generically non-differentiable (Laurent & Brecht, 2018). As a result, the points of interest lie where their approximation breaks down.
To challenge the broad applicability of their statement, consider the Huber loss function depicted in Equation (13). This function is overparameterized and therefore possesses a global minimum of 0. Consequently, within the vicinity of this global minimum, the loss function exhibits quadratic growth. In such instances, Ma et al. (2022)’s result states that we will not see unstable convergence. However, contrary to their suggestions, Proposition 4 guarantees unstable convergence under these circumstances, as evidenced in Figure 3.
The discrepancy arises because the Huber loss function lacks the twice differentiability required by Definition 4.1 of Ma et al. (2022)’s paper. In contrast, all $L-smooth$ functions will exhibit subquadratic growth near their local minimum during unstable convergence, as noted by (Ma et al., 2022). This occurs because $L-smooth$ functions are twice differentiable by definition. This example highlights the challenges of extending results developed under strong assumptions like $L-smoothness$ to the general case.

### v3

@@ 3 NDGM and the LASSO penalty
Zero weights are crucial for model compression (Blalock et al., 2020; Han et al., 2015a, b; Li et al., 2016), and inducing sparsity by penalizing the $L_{1}$ norm of model weights is a common approach. While the statistics literature employs specialized algorithms for problems involving penalized $L_{1}$ norms (Efron et al., 2004; Friedman et al., 2008; Mazumder & Hastie, 2012; Tibshirani, 1996; Tibshirani et al., 2005; Zou & Hastie, 2005), neural networks often use the penalized $L_{1}$ norm with arbitrary non-convex functions, assuming that NDGM results in a near-sparse optimal solution (Bengio, 2012; Goodfellow et al., 2016b; Scardapane et al., 2017). This expectation is not always met, as evidenced by user struggles in PyTorch forums [^4] https://rb.gy/kh0qr [^5] https://rb.gy/2vsry [^6] https://rb.gy/6dfli. Given the LASSO’s critical applications (Ghosh et al., 2021; Li et al., 2011; Li & Sillanpää, 2012; Ogutu et al., 2012), understanding the reliability of the NDGM solution for the LASSO problem is essential.
@@
Let $\{\beta_{t}\}_{t=0}^{\infty}$ denote the sequence generated by running NDGM on (7), with a constant learning rate, $\alpha$, and the entries in $\beta_{0}$ being uniformly sampled from $[-1,1]$. Then $\beta_{k}$ will have all non-zero entries with probability 1 for all $k>0$. Furthermore, there exists an integer $N_{0}$, such that for every $N\geq N_{0}$, we can write
$\beta_{n_{k}+m}[k]=\begin{cases}\gamma_{k}-\alpha\lambda_{1}&\text{ if m is odd}\\ \gamma_{k}&\text{ if m is even},\end{cases}$ (8)
where $\gamma_{k}\in(0,\alpha\lambda_{1})$ for all $10\leq k\leq P-$ 1.
@@
See Appendix B for details ∎
Equation (8) suggests that with the same learning rate and initialization, a larger $\lambda_{1}$ can lead to a NDGM solution with a larger $L_{1}$ norm, defeating the purpose of the LASSO penalty. To illustrate this, we note that in the 2D case, with weights initialized as $\beta_{0}=[0.5053,0.5053]$, $\alpha=0.01$, and $\lambda_{1}=1$, the NDGM sequence oscillates between $[0.0053,0.0053]$ and $[-0.0047,-0.0047]$. However, with $\lambda_{1}=100$, and the same values of $\beta_{0}$ and $\alpha_{1}$, the solution oscillates between $[0.5053,0.5053]$ and $[-0.4947,-0.4947]$. Therefore at convergence, the $L_{1}$ norm is at most 0.0106 for $\lambda_{1}=1$, but at least 0.9894 for $\lambda_{1}=100$. Furthermore, reducing the step-size to $\alpha=0.001$ after apparent convergence does not resolve this paradox. With $\lambda_{1}=1$, the sequence oscillates between $[0.0003,0.0003]$ and $[-0.0007,-0.0007]$, while with $\lambda_{1}=100$, it oscillates between $[0.0053,0.0053]$ and $[-0.0947,-0.0947]$. Appendix C shows that this paradox holds for NDGM solutions of the standard $L_{1}$ regularized linear problem.
The subgradient method and step-size choices:
The counter-intuitive behavior seen in the example above can be explained using the subgradient method. Since the loss function described in equation (7) is Lipschitz continuous (equation (6)) with $L=\lambda_{1}$, we have from (Boyd et al., 2003) (Section 2 on the Constant Step Size) that
@@
Therefore, for the same value of $\alpha$, the larger value of $\lambda_{1}$ will result in a larger error in estimation, which is precisely what we are seeing above. This illustration also highlights the contrast between the diminishing and reducing step size regimes outlined in the preliminaries. Under the diminishing step-size regime, $\alpha\rightarrow 0$ in the theoretical limit, and therefore, Proposition 3.1 guarantees that the optimal solution for (7) will be 0 for all $\lambda_{1}>0$. On the other hand, in the reducing step-size regime, $\alpha>0$ at the end of training, and therefore, the error in estimation becomes a critical determinant of the optimal value.
Connections to other NDGM variants:
Conventional wisdom in deep learning theory suggests that different NDGM variants, such as NDGM with momentum and RMSProp, converge at varying rates but eventually reach similar near-optimal values (Figure 8.5 of (Goodfellow et al., 2016a), slides 64 onwards in https://rb.gy/ojzpgm  [^7] These slides are from one of the most widely used courses for self teaching; see 1:01:00 - 1:03:00 of https://rb.gy/kaj976 where this precise point is made). However, we prove in Appendices D and E that these claims are incorrect even for the toy LASSO problem described in (7). Furthermore, our analysis reveals two interesting facts. First, that the RMSProp sequence for (7) is agnostic of the value of $\lambda_{1}$ used in the analysis. Second, there is a phase transition between the RMSProp and vanilla NDGM training curves characterized by a change in the relative behavior of the training curves of the two algorithms as $\lambda_{1}$ is increased from a value which is much smaller than 1 to a value much larger than 1, which cannot be explained without accounting for the non-differentiability of the $L_{1}$ regularizer. These findings highlight the importance of incorporating the non-differentiability of regularizers like the $L_{1}$ norm when analyzing the behavior of NDGM variants. Failing to account for non-differentiability results in an incomplete understanding of the optimization dynamics. Consequently, it remains unclear how the convergence analysis of algorithms like Adam and RMSProp for differentiable loss functions, as often performed in studies such as (Reddi et al., 2019; Défossez et al., 2020; Zhang et al., 2022), relates to the dynamics of these algorithms’ convergence in non-differentiable settings involving regularizers like the $L_{1}$ norm or activation functions like the ReLU.
@@
In their seminal work on network pruning, (Han et al., 2015b) propose using $L_{1}$ regularization to encourage weights to approach 0, facilitating the pruning process. This approach is widely adopted in the literature (Wen et al., 2016; Xiao et al., 2019; Yang et al., 2019) and industry [^8] See the recommendation of the regularizer in the Training config section of https://rb.gy/20y13m and https://rb.gy/i1sb6s. To challenge this suggestion’s general applicability, we conduct two sets of experiments finetuning a pre-trained VGG16 model on CIFAR-10 using $L_{1}$ penalized cross-entropy loss with a constant learning rate of $\alpha=0.003$ for 200 epochs. We use $\lambda_{1}=0.0002$ in the first set and $\lambda_{1}=2$ in the second, running three parallel trials within each set using RMSProp, SGD, and Adam. At the end of our runs, we make two key observations (Figure 5): 1) The $L_{1}$ norms at convergence differ vastly across the three variants, and 2) for SGD and Adam, larger values of $\lambda_{1}$ result in a larger $L_{1}$ norm at convergence. These findings suggest that the relationship between $L_{1}$ regularization and prunable parameters may not be as straightforward as previously thought. Indeed, our results show that a larger $\lambda_{1}$ does not necessarily correspond to more prunable parameters (Table 1): with $\lambda_{1}=2$, less than 0.2% of the parameters at convergence with vanilla NDGM have absolute value less than $10^{-5}$, but with $\lambda_{1}=0.002$, 99.98% do.
Soft-thresholding and ISTA:
The observation that higher $\lambda_{1}$ values do not yield more prunable parameters indicates that while soft thresholding produces sparse solutions for NDGMs, it does not find optimal solutions. This issue extends beyond NDGM to general neural network problems, where any soft thresholding solution yields sparsity but not optimality. The root cause is that soft thresholding implicitly solves a generalized ISTA problem (Chen et al., 2023), which assumes the loss function is $L-smooth$ (point 1, Section 3.2 of (Chen et al., 2023)) - an assumption that fails to capture the non-convex and non-differentiable dynamics in general neural network settings, as discussed previously. Consequently, applying the generalized ISTA formulation relying on $L-smoothness$ to neural networks necessarily leads to suboptimal solutions.
@@
While NDGMs do not inherently generate sparse solutions, iterative methods have been developed for $L_{1}$ regularized problems (Ma et al., 2019; Siegel et al., 2020). The realization that NDGMs lack this capability predates the widespread adoption of deep learning (Xiao, 2009). Despite this, a misconception persists in the deep learning literature, suggesting that such algorithms yield sparse solutions comparable to those from traditional machine learning frameworks like glmnet (Friedman et al., 2021) or scikit-learn (Kramer & Kramer, 2016). However, there is a fundamental disparity between these frameworks: glmnet focuses on a narrow set of problems, guaranteeing optimality, while PyTorch (Paszke et al., 2017) provides flexibility in model creation, but lacks similar guarantees. Attaining a solution at a non-differentiable point is challenging, emphasizing the importance of exploring related studies and tools rather than relying on simplistic implementations when precision is crucial.
@@ 4 NDGM and the Edge of Stability
In recent years, there has been a growing interest on the topic of “unstable convergence”. The literature on the subject (Ahn et al., 2022b, a; Arora et al., 2022; Chen & Bruna, 2022; Cohen et al., 2021; Li et al., 2022) is motivated by the finding that unlike convex quadratic forms, gradient descent on neural networks do not diverge to $\infty$ even when the learning rate is greater than $\alpha^{*}=2/\eta$, where $\eta$ is the dominant eigenvalue of the loss function’s Hessian. Instead, the loss function has been empirically demonstrated to converge “unstably”, reducing non-montonically on the long run. $\alpha=\alpha^{*}$ is referred to in the literature as the ”Edge of Stability”, since it separates regions of ”stable convergence” ($\alpha<\alpha^{*}$) from regions of “unstable convergence” [^9] We define unstable convergence to mean “non-divergence to $\infty$”. Our definition is adapted from Definition 1.1 of (Arora et al., 2022). However, we deviate from their definition due to its dependence on the existence of the Hessian at all points between the current and next iterate, a condition that need not be satisfied in non-smooth neural networks. ($\alpha>\alpha^{*}$). The derivation of the Edge of Stability condition assumes that the loss function, $f$, is $L-smooth$ (see equation (5)).
It is conjectured that neural networks demonstrate unstable convergence because the phenomenon is inconsistent with widespread presumptions in the field of optimization (Cohen et al., 2021; Ahn et al., 2024). This conjecture is incorrect for Lipschitz continuous, differentiable, convex functions satisfying (6) as shown in the proposition below.
@@
The toy LASSO problem studied in (7), for which we analytically demonstrated unstable convergence in the previous section, is an example of a non-differentiable convex loss function satisfying Proposition 4.1. An example of a differentiable non-smooth convex function satisfying Proposition 4.1 is the Huber loss function, which finds use in many state of the art object detection models (Girshick, 2015; Liu et al., 2016; Ren et al., 2015). The Huber loss function is non-smooth because it is once but not twice differentiable. For a regression problem involving an arbitrary $50\times 1$ response vector, $\mathbf{y}=[y_{1},y_{2}\ldots y_{N}]^{T}$, and an arbitrary $50\times 200$ data matrix, $\mathbf{Z}$, the Huber loss function is defined as
@@
where $z_{i}$ is the $i^{th}$ column of $\mathbf{Z}$. Figure 3 shows that the NDGM sequence for (10) does not diverge toward $\infty$, even with a high learning rate, such as $\alpha=10$, which aligns with our expectations. Proposition 4.1 suggest that it is unclear and the examples that follow show that the common approach of extrapolating results developed for $L-smooth$ loss functions to the non-smooth or non-differentiable case is problematic. To emphasize this point, we provide an example of such an extrapolation in the literature which breaks down in the non-smooth case.
Subquadratic losses and unstable convergence:
In section 3.1 of (Ma et al., 2022), the authors use second-order finite differences to analyze the curvature of VGG16 and ResNet near the local minimum, concluding that subquadratic growth explains the edge of stability phenomenon. However, their analysis is flawed because ReLU networks have generically non-differentiable local minima (Laurent & Brecht, 2018), where their approximation breaks down.
The Huber loss function described in equation (10) challenges their statement’s broad applicability. This overparameterized convex function has a global minimum of 0, exhibiting quadratic growth nearby. (Ma et al., 2022)’s result suggests stable convergence, but Proposition 4 guarantees unstable convergence (Figure 3). The discrepancy arises because the Huber loss function lacks the twice differentiability required by Definition 4.1 in (Ma et al., 2022). In contrast, $L-smooth$ functions, which are twice differentiable by definition, exhibit subquadratic growth near their local minimum during unstable convergence. This example highlights the challenges of extending results developed under strong assumptions like $L-smoothness$ to the general case.